$SPAD/input schaum31.input

Timothy Daly December 3, 2016

Contents

2

1 [?]:14.615 ∫ coth ax dx

∫ coth ax = \frac{1}{a} ln sinh ax

— * —

)set break resume
)spool schaum31.output
)set message test on
)set message auto off
)clear all
--S 1 of 46
aa:=integrate(coth(a*x),x)
--R
--R
--R                    2sinh(a x)
--R        log(- ---------------------) - a x
--R              sinh(a x) - cosh(a x)
--R   (1)  ----------------------------------
--R                         a
--R                                         Type: Union(Expression(Integer),...)
--E
--S 2 of 46
bb:=1/a*log(sinh(a*x))
--R
--R        log(sinh(a x))
--R   (2)  --------------
--R               a
--R                                                    Type: Expression(Integer)
--E
--S 3 of 46
cc:=aa-bb
--R
--R                                       2sinh(a x)
--R        - log(sinh(a x)) + log(- ---------------------) - a x
--R                                 sinh(a x) - cosh(a x)
--R   (3)  -----------------------------------------------------
--R                                  a
--R                                                    Type: Expression(Integer)
--E
--S 4 of 46
dd:=expandLog cc
--R
--R        - log(sinh(a x) - cosh(a x)) + log(- 2) - a x

--R   (4)  ---------------------------------------------
--R                              a
--R                                                    Type: Expression(Integer)
--E
--S 5 of 46      14:615 Schaums and Axiom differ by a constant
ee:=complexNormalize dd
--R
--R        - log(- 1) + log(- 2)
--R   (5)  ---------------------
--R                  a
--R                                                    Type: Expression(Integer)
--E

———-

2 [?]:14.616 ∫ coth^{2} ax dx

∫ coth^{2} ax = x − \frac{coth ax}{a}

— * —

)clear all
--S 6 of 46
aa:=integrate(coth(a*x)^2,x)
--R
--R
--R        (a x + 1)sinh(a x) - cosh(a x)
--R   (1)  ------------------------------
--R                  a sinh(a x)
--R                                         Type: Union(Expression(Integer),...)
--E
--S 7 of 46
bb:=x-coth(a*x)/a
--R
--R        - coth(a x) + a x
--R   (2)  -----------------
--R                a
--R                                                    Type: Expression(Integer)
--E
--S 8 of 46
cc:=aa-bb
--R
--R        (coth(a x) + 1)sinh(a x) - cosh(a x)
--R   (3)  ------------------------------------

--R                     a sinh(a x)
--R                                                    Type: Expression(Integer)
--E
--S 9 of 46      14:616 Schaums and Axiom differ by a constant
dd:=complexNormalize cc
--R
--R        1
--R   (4)  -
--R        a
--R                                                    Type: Expression(Integer)
--E

———-

3 [?]:14.617 ∫ coth^{3} ax dx

∫ coth^{3} ax = \frac{1}{a} ln sinh ax − \frac{coth^{2} ax}{2a}

— * —

)clear all
--S 10 of 46
aa:=integrate(coth(a*x)^3,x)
--R
--R
--R   (1)
--R                    4                      3              2              2
--R           sinh(a x)  + 4cosh(a x)sinh(a x)  + (6cosh(a x)  - 2)sinh(a x)
--R         +
--R                      3                                   4             2
--R           (4cosh(a x)  - 4cosh(a x))sinh(a x) + cosh(a x)  - 2cosh(a x)  + 1
--R      *
--R                     2sinh(a x)
--R         log(- ---------------------)
--R               sinh(a x) - cosh(a x)
--R     +
--R                      4                          3
--R       - a x sinh(a x)  - 4a x cosh(a x)sinh(a x)
--R     +
--R                        2                     2
--R       (- 6a x cosh(a x)  + 2a x - 2)sinh(a x)
--R     +
--R                        3                                                4
--R       (- 4a x cosh(a x)  + (4a x - 4)cosh(a x))sinh(a x) - a x cosh(a x)
--R     +
--R                          2

--R       (2a x - 2)cosh(a x)  - a x
--R  /
--R                  4                        3                2               2
--R       a sinh(a x)  + 4a cosh(a x)sinh(a x)  + (6a cosh(a x)  - 2a)sinh(a x)
--R     +
--R                  3                                       4               2
--R     (4a cosh(a x)  - 4a cosh(a x))sinh(a x) + a cosh(a x)  - 2a cosh(a x)  + a
--R                                         Type: Union(Expression(Integer),...)
--E
--S 11 of 46
bb:=1/a*log(sinh(a*x)-coth(a*x)^2)/(2*a)
--R
--R                                 2
--R        log(sinh(a x) - coth(a x) )
--R   (2)  ---------------------------
--R                      2
--R                    2a
--R                                                    Type: Expression(Integer)
--E
--S 12 of 46     14:617 Axiom cannot simplify this expression
cc:=aa-bb
--R
--R   (3)
--R                      4                      3                2              2
--R           - sinh(a x)  - 4cosh(a x)sinh(a x)  + (- 6cosh(a x)  + 2)sinh(a x)
--R         +
--R                        3                                   4             2
--R           (- 4cosh(a x)  + 4cosh(a x))sinh(a x) - cosh(a x)  + 2cosh(a x)  - 1
--R      *
--R                                  2
--R         log(sinh(a x) - coth(a x) )
--R     +
--R                       4                        3
--R           2a sinh(a x)  + 8a cosh(a x)sinh(a x)
--R         +
--R                         2               2
--R           (12a cosh(a x)  - 4a)sinh(a x)
--R         +
--R                        3                                        4
--R           (8a cosh(a x)  - 8a cosh(a x))sinh(a x) + 2a cosh(a x)
--R         +
--R                         2
--R           - 4a cosh(a x)  + 2a
--R      *
--R                     2sinh(a x)
--R         log(- ---------------------)
--R               sinh(a x) - cosh(a x)
--R     +

--R           2           4     2                    3
--R       - 2a x sinh(a x)  - 8a x cosh(a x)sinh(a x)
--R     +
--R             2           2     2                2
--R       (- 12a x cosh(a x)  + 4a x - 4a)sinh(a x)
--R     +
--R            2           3      2                               2           4
--R       (- 8a x cosh(a x)  + (8a x - 8a)cosh(a x))sinh(a x) - 2a x cosh(a x)
--R     +
--R          2                2     2
--R       (4a x - 4a)cosh(a x)  - 2a x
--R  /
--R         2         4     2                  3       2         2     2          2
--R       2a sinh(a x)  + 8a cosh(a x)sinh(a x)  + (12a cosh(a x)  - 4a )sinh(a x)
--R     +
--R          2         3     2                        2         4     2         2
--R       (8a cosh(a x)  - 8a cosh(a x))sinh(a x) + 2a cosh(a x)  - 4a cosh(a x)
--R     +
--R         2
--R       2a
--R                                                    Type: Expression(Integer)
--E

———-

4 [?]:14.618 ∫ coth^{n} ax csch^{2}ax dx

∫ coth^{n} ax csch^{2}ax = −\frac{coth^{n+1} ax}{(n + 1)a}

— * —

)clear all
--S 13 of 46
aa:=integrate(coth(a*x)^n*csch(a*x)^2,x)
--R
--R
--R                              cosh(a x)                         cosh(a x)
--R        - cosh(a x)sinh(n log(---------)) - cosh(a x)cosh(n log(---------))
--R                              sinh(a x)                         sinh(a x)
--R   (1)  -------------------------------------------------------------------
--R                                 (a n + a)sinh(a x)
--R                                         Type: Union(Expression(Integer),...)
--E
--S 14 of 46
bb:=-coth(a*x)^(n+1)/((n+1)*a)
--R

--R                   n + 1
--R          coth(a x)
--R   (2)  - --------------
--R              a n + a
--R                                                    Type: Expression(Integer)
--E
--S 15 of 46
cc:=aa-bb
--R
--R   (3)
--R                             cosh(a x)                         cosh(a x)
--R       - cosh(a x)sinh(n log(---------)) - cosh(a x)cosh(n log(---------))
--R                             sinh(a x)                         sinh(a x)
--R     +
--R                         n + 1
--R       sinh(a x)coth(a x)
--R  /
--R     (a n + a)sinh(a x)
--R                                                    Type: Expression(Integer)
--E
--S 16 of 46
dd:=expandLog cc
--R
--R   (4)
--R       cosh(a x)sinh(n log(sinh(a x)) - n log(cosh(a x)))
--R     +
--R       - cosh(a x)cosh(n log(sinh(a x)) - n log(cosh(a x)))
--R     +
--R                         n + 1
--R       sinh(a x)coth(a x)
--R  /
--R     (a n + a)sinh(a x)
--R                                                    Type: Expression(Integer)
--E
--S 17 of 46     14:618 Schaums and Axiom agree
ee:=complexNormalize dd
--R
--R   (5)  0
--R                                                    Type: Expression(Integer)
--E

———-

5 [?]:14.619 ∫ \begin{matrix}csch & 2 & ax \\ coth & ax \end{matrix} dx

∫ \begin{matrix}csch & 2 & ax \\ coth & ax \end{matrix} = −\frac{1}{a} ln coth ax

— * —

)clear all
--S 18 of 46
aa:=integrate(csch(a*x)^2/coth(a*x),x)
--R
--R
--R                      2cosh(a x)                     2sinh(a x)
--R        - log(- ---------------------) + log(- ---------------------)
--R                sinh(a x) - cosh(a x)          sinh(a x) - cosh(a x)
--R   (1)  -------------------------------------------------------------
--R                                      a
--R                                         Type: Union(Expression(Integer),...)
--E
--S 19 of 46
bb:=-1/a*log(coth(a*x))
--R
--R          log(coth(a x))
--R   (2)  - --------------
--R                 a
--R                                                    Type: Expression(Integer)
--E
--S 20 of 46
cc:=aa-bb
--R
--R   (3)
--R                                2cosh(a x)                     2sinh(a x)
--R   log(coth(a x)) - log(- ---------------------) + log(- ---------------------)
--R                          sinh(a x) - cosh(a x)          sinh(a x) - cosh(a x)
--R   ----------------------------------------------------------------------------
--R                                         a
--R                                                    Type: Expression(Integer)
--E
--S 21 of 46
dd:=expandLog cc
--R
--R        log(sinh(a x)) + log(coth(a x)) - log(cosh(a x))
--R   (4)  ------------------------------------------------
--R                                a

--R                                                    Type: Expression(Integer)
--E
--S 22 of 46     14:619 Schaums and Axiom agree
ee:=complexNormalize dd
--R
--R   (5)  0
--R                                                    Type: Expression(Integer)
--E

———-

6 [?]:14.620 ∫ \frac{dx}{coth ax} dx

∫ \frac{1}{coth ax} = \frac{1}{a} ln cosh ax

— * —

)clear all
--S 23 of 46
aa:=integrate(1/coth(a*x),x)
--R
--R
--R                    2cosh(a x)
--R        log(- ---------------------) - a x
--R              sinh(a x) - cosh(a x)
--R   (1)  ----------------------------------
--R                         a
--R                                         Type: Union(Expression(Integer),...)
--E
--S 24 of 46
bb:=1/a*log(cosh(a*x))
--R
--R        log(cosh(a x))
--R   (2)  --------------
--R               a
--R                                                    Type: Expression(Integer)
--E
--S 25 of 46
cc:=aa-bb
--R
--R                                       2cosh(a x)
--R        - log(cosh(a x)) + log(- ---------------------) - a x
--R                                 sinh(a x) - cosh(a x)

--R   (3)  -----------------------------------------------------
--R                                  a
--R                                                    Type: Expression(Integer)
--E
--S 26 of 46
dd:=expandLog cc
--R
--R        - log(sinh(a x) - cosh(a x)) + log(- 2) - a x
--R   (4)  ---------------------------------------------
--R                              a
--R                                                    Type: Expression(Integer)
--E
--S 27 of 46     14:620 Schaums and Axiom differ by a constant
ee:=complexNormalize dd
--R
--R        - log(- 1) + log(- 2)
--R   (5)  ---------------------
--R                  a
--R                                                    Type: Expression(Integer)
--E

———-

7 [?]:14.621 ∫ x coth ax dx

∫ x coth ax = \frac{1}{a^{2}} \{ax + \frac{(ax)^{3}}{9} − \frac{(ax)^{5}}{225} + ⋯ \frac{(−1)^{n−1}2^{2n}B_{n}(ax)^{2n+1}}{(2n + 1)!} + ⋯\}

— * —

)clear all
--S 28 of 46     14:621 Axiom cannot compute this integral
aa:=integrate(x*coth(a*x),x)
--R
--R
--R           x
--R         ++
--I   (1)   |   %O coth(%O a)d%O
--R        ++
--R                                         Type: Union(Expression(Integer),...)
--E

———-

8 [?]:14.622 ∫ x coth^{2} ax dx

∫ x coth^{2} ax = \frac{x^{2}}{2} − \frac{x coth ax}{a} + \frac{1}{a^{2}} ln sinh ax

— * —

)clear all
--S 29 of 46
aa:=integrate(x*coth(a*x)^2,x)
--R
--R
--R   (1)
--R                    2                                   2
--R         (2sinh(a x)  + 4cosh(a x)sinh(a x) + 2cosh(a x)  - 2)
--R      *
--R                     2sinh(a x)
--R         log(- ---------------------)
--R               sinh(a x) - cosh(a x)
--R     +
--R         2 2                 2      2 2
--R       (a x  - 4a x)sinh(a x)  + (2a x  - 8a x)cosh(a x)sinh(a x)
--R     +
--R         2 2                 2    2 2
--R       (a x  - 4a x)cosh(a x)  - a x
--R  /
--R       2         2     2                       2         2     2
--R     2a sinh(a x)  + 4a cosh(a x)sinh(a x) + 2a cosh(a x)  - 2a
--R                                         Type: Union(Expression(Integer),...)
--E
--S 30 of 46
bb:=x^2/2-(x*coth(a*x)/a)+1/a^2*log(sinh(a*x))
--R
--R                                            2 2
--R        2log(sinh(a x)) - 2a x coth(a x) + a x
--R   (2)  ---------------------------------------
--R                            2
--R                          2a
--R                                                    Type: Expression(Integer)
--E
--S 31 of 46
cc:=aa-bb
--R
--R   (3)
--R                   2                                  2
--R       (- sinh(a x)  - 2cosh(a x)sinh(a x) - cosh(a x)  + 1)log(sinh(a x))

--R     +
--R                   2                                  2
--R         (sinh(a x)  + 2cosh(a x)sinh(a x) + cosh(a x)  - 1)
--R      *
--R                     2sinh(a x)
--R         log(- ---------------------)
--R               sinh(a x) - cosh(a x)
--R     +
--R                                      2
--R       (a x coth(a x) - 2a x)sinh(a x)
--R     +
--R       (2a x cosh(a x)coth(a x) - 4a x cosh(a x))sinh(a x)
--R     +
--R                     2                                 2
--R       (a x cosh(a x)  - a x)coth(a x) - 2a x cosh(a x)
--R  /
--R      2         2     2                      2         2    2
--R     a sinh(a x)  + 2a cosh(a x)sinh(a x) + a cosh(a x)  - a
--R                                                    Type: Expression(Integer)
--E
--S 32 of 46
dd:=expandLog cc
--R
--R   (4)
--R                     2                                  2
--R         (- sinh(a x)  - 2cosh(a x)sinh(a x) - cosh(a x)  + 1)
--R      *
--R         log(sinh(a x) - cosh(a x))
--R     +
--R                                                 2
--R       (a x coth(a x) + log(- 2) - 2a x)sinh(a x)
--R     +
--R       (2a x cosh(a x)coth(a x) + (2log(- 2) - 4a x)cosh(a x))sinh(a x)
--R     +
--R                     2                                             2
--R       (a x cosh(a x)  - a x)coth(a x) + (log(- 2) - 2a x)cosh(a x)  - log(- 2)
--R  /
--R      2         2     2                      2         2    2
--R     a sinh(a x)  + 2a cosh(a x)sinh(a x) + a cosh(a x)  - a
--R                                                    Type: Expression(Integer)
--E
--S 33 of 46
sinhsqrrule:=rule(sinh(x)^2 == 1/2*cosh(2*x)-1/2)
--R
--R               2    cosh(2x) - 1
--R   (5)  sinh(x)  == ------------
--R                          2
--R                       Type: RewriteRule(Integer,Integer,Expression(Integer))

--E
--S 34 of 46
ee:=sinhsqrrule dd
--R
--R   (6)
--R                                                         2
--R         (- 4cosh(a x)sinh(a x) - cosh(2a x) - 2cosh(a x)  + 3)
--R      *
--R         log(sinh(a x) - cosh(a x))
--R     +
--R       (4a x cosh(a x)coth(a x) + (4log(- 2) - 8a x)cosh(a x))sinh(a x)
--R     +
--R                                       2
--R       (a x cosh(2a x) + 2a x cosh(a x)  - 3a x)coth(a x)
--R     +
--R                                                                2
--R       (log(- 2) - 2a x)cosh(2a x) + (2log(- 2) - 4a x)cosh(a x)  - 3log(- 2)
--R     +
--R       2a x
--R  /
--R       2                      2               2         2     2
--R     4a cosh(a x)sinh(a x) + a cosh(2a x) + 2a cosh(a x)  - 3a
--R                                                    Type: Expression(Integer)
--E
--S 35 of 46
coshsqrrule:=rule(cosh(x)^2 == 1/2*cosh(2*x)+1/2)
--R
--R               2    cosh(2x) + 1
--R   (7)  cosh(x)  == ------------
--R                          2
--R                       Type: RewriteRule(Integer,Integer,Expression(Integer))
--E
--S 36 of 46
ff:=coshsqrrule ee
--R
--R   (8)
--R       (- 2cosh(a x)sinh(a x) - cosh(2a x) + 1)log(sinh(a x) - cosh(a x))
--R     +
--R       (2a x cosh(a x)coth(a x) + (2log(- 2) - 4a x)cosh(a x))sinh(a x)
--R     +
--R       (a x cosh(2a x) - a x)coth(a x) + (log(- 2) - 2a x)cosh(2a x) - log(- 2)
--R  /
--R       2                      2              2
--R     2a cosh(a x)sinh(a x) + a cosh(2a x) - a
--R                                                    Type: Expression(Integer)
--E

--S 37 of 46
sinhcoshrule:=rule(sinh(x)*cosh(y) == 1/2*(sinh(x+y)+sinh(x-y)))
--R
--I                             %L sinh(y + x) - %L sinh(y - x)
--I   (9)  %L cosh(y)sinh(x) == -------------------------------
--R                                            2
--R                       Type: RewriteRule(Integer,Integer,Expression(Integer))
--E
--S 38 of 46
gg:=sinhcoshrule ff
--R
--R   (10)
--R       (- sinh(2a x) - cosh(2a x) + 1)log(sinh(a x) - cosh(a x))
--R     +
--R       (a x coth(a x) + log(- 2) - 2a x)sinh(2a x)
--R     +
--R       (a x cosh(2a x) - a x)coth(a x) + (log(- 2) - 2a x)cosh(2a x) - log(- 2)
--R  /
--R      2              2              2
--R     a sinh(2a x) + a cosh(2a x) - a
--R                                                    Type: Expression(Integer)
--E
--S 39 of 46     14:622 Schaums and Axiom differ by a constant
hh:=complexNormalize gg
--R
--R         - log(- 1) + log(- 2)
--R   (11)  ---------------------
--R                    2
--R                   a
--R                                                    Type: Expression(Integer)
--E

———-

9 [?]:14.623 ∫ \frac{coth ax}{x} dx

∫ \frac{coth ax}{x} = −\frac{1}{ax} + \frac{(ax)}{3} − \frac{(ax)^{3}}{135} + ⋯ \frac{(−1)^{n}2^{2n}B_{n}(ax)^{2n−1}}{(2n − 1)(2n)!} + ⋯

— * —

)clear all
--S 40 of 46     14:623 Axiom cannot compute this integral
aa:=integrate(coth(a*x)/x,x)
--R

--R
--R           x
--I         ++  coth(%O a)
--I   (1)   |   ---------- d%O
--I        ++       %O
--R                                         Type: Union(Expression(Integer),...)
--E

———-

10 [?]:14.624 ∫ \frac{dx}{p + q coth ax} dx

∫ \frac{1}{p + q coth ax} = \frac{px}{p^{2} − q^{2}} − \frac{q}{a(p^{2} − q^{2})} ln(p sinh ax + q cosh ax)

— * —

)clear all
--S 41 of 46
aa:=integrate(1/(p+q*coth(a*x)),x)
--R
--R
--R              - 2p sinh(a x) - 2q cosh(a x)
--R        q log(-----------------------------) + (- a q - a p)x
--R                  sinh(a x) - cosh(a x)
--R   (1)  -----------------------------------------------------
--R                                2      2
--R                             a q  - a p
--R                                         Type: Union(Expression(Integer),...)
--E
--S 42 of 46
bb:=(p*x)/(p^2-q^2)-q/(a*(p^2-q^2))*log(p*sinh(a*x)+q*cosh(a*x))
--R
--R        q log(p sinh(a x) + q cosh(a x)) - a p x
--R   (2)  ----------------------------------------
--R                          2      2
--R                       a q  - a p
--R                                                    Type: Expression(Integer)
--E
--S 43 of 46
cc:=aa-bb
--R
--R   (3)
--R                                                  - 2p sinh(a x) - 2q cosh(a x)
--R       - q log(p sinh(a x) + q cosh(a x)) + q log(-----------------------------)

--R                                                      sinh(a x) - cosh(a x)
--R     +
--R       - a q x
--R  /
--R        2      2
--R     a q  - a p
--R                                                    Type: Expression(Integer)
--E
--S 44 of 46
dd:=expandLog cc
--R
--R   (4)
--R       - q log(p sinh(a x) + q cosh(a x)) - q log(sinh(a x) - cosh(a x))
--R     +
--R       q log(- p sinh(a x) - q cosh(a x)) + q log(2) - a q x
--R  /
--R        2      2
--R     a q  - a p
--R                                                    Type: Expression(Integer)
--E
--S 45 of 46     14:624 Schaums and Axiom differ by a constant
ee:=complexNormalize dd
--R
--R        q log(2) - 2q log(- 1)
--R   (5)  ----------------------
--R                 2      2
--R              a q  - a p
--R                                                    Type: Expression(Integer)
--E

———-

11 [?]:14.625 ∫ coth^{n} ax dx

∫ coth^{n} ax = −\frac{coth^{n−1} ax}{a(n − 1)} + ∫ coth^{n−1} ax

— * —

)clear all
--S 46 of 46     14:625 Axiom cannot compute this integral
aa:=integrate(coth(a*x)^n,x)
--R
--R
--R           x
--R         ++            n

--I   (1)   |   coth(%O a) d%O
--R        ++
--R                                         Type: Union(Expression(Integer),...)
--E
)spool
)lisp (bye)

———-

References

[1] Spiegel, Murray R. Mathematical Handbook of Formulas and Tables Schaum’s Outline Series McGraw-Hill 1968 pp90-91